If A and B are mutually exclusive, with P(A) = 0.33, and P(B) = 0.28, find

(a)P(Ac),

(b)P(Bc),

(c)P(A∪B),

(d)P(A∩B),

(e)P(Ac∩B),

(f)P(Ac|Bc).
Answer #1

Solution:

If A and B are mutually exclusive, with

P(A) = 0.33, and P(B) = 0.28,

(a)P(A^c)=1-P(A)=1-0.33=0.67

(b)P(B^c)=1-P(B)=1-0.28=0.72

(c)P(A \cup B)= P(A)+P(B)-P(A \cap B)

=P(A)+P(B)=0.33+0.28=0.61

(d)P(A \cap B)=0,  As A and B are mutually exclusive.

(e)P(A^c \cap B) = P(B)- P(A \cap B)= 0.28- 0 =0.28

(f)P(A^c|B^c)=\frac{P(A^c \cap B^c)}{P(B^c)}

P(A^c|B^c)=\frac{1-P(A \cup B)^c}{P(B^c)}

P(A^c|B^c)=\frac{1-0.61}{0.72}

P(A^c|B^c)=\frac{0.39}{0.72} = 0.54166= 0.54

**For further queries, please comment below.

Thank you.